OpenStudy (anonymous):

The city planning committee is composed of 12 women and 14 men. What fraction of the committee is composed of women? A. 1/12 B. 6/13 C. 7/13 D. 6/7

OpenStudy (kira_yamato):

12 / (12 + 14) = 12/26 = 6/13